# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ Appendix A Properties of the Multinomial Logit Function
Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}\leq 1$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
@@
$\sum_{i=1}^{k}\sum_{s^{\prime}\in\dot{\mathcal{S}}_{i,h}}p_{i,h}^{s^{\prime}}(\widetilde{\theta}_{i+1,h})p_{i,h}^{\dot{s}_{i,h}}(\widetilde{\theta}_{i+1,h})\lVert\phi_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}}$
@@
Proof of statement (III).  By Lemma 5, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k+1,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$
Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (2), we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h})}{\det(\mathcal{H}_{1,h})}\right)$
$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
This finishes the proof.
Proof of statement (IV).  By the definition of $H_{k,h}(\theta)$, we have
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}(\phi_{i,h}^{s^{\prime}})^{\top}]-\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{(}\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}[\phi_{i,h}^{s^{\prime}}]\big{)}^{\top}$
$=\mathbb{E}_{s^{\prime}\in p_{i,h}(\theta)}\big{[}(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})(\phi_{i,h}^{s^{\prime}}-\mathbb{E}_{s^{\prime\prime}\in p_{i,h}(\theta)}\phi_{i,h}^{s^{\prime\prime}})^{\top}\big{]}$
Thus, we have
$H_{i,h}(\widetilde{\theta}_{i+1,h})\succeq\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\widetilde{\theta}_{i+1,h})(\widetilde{\phi}_{i,h}^{s^{\prime}})(\widetilde{\phi}_{i,h}^{s^{\prime}})^{\top}.$ (13)
Then, we get
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}(\widetilde{\theta}_{k+1,h})(\widetilde{\phi}_{k,h}^{s^{\prime}})(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}$
The remaining proof is similar to the proof of statement (2).
Proof of statement (V).  By (13), we have
Then, the remaining proof is similar to the proof of statement (3). ∎
Appendix B Proof of Lemma 1
B.1 Main Proof
Proof.
The main proof follows the proof of Theorem 1 in Zhang et al. [2016]. We define
$\bar{f}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[f_{k,h}(\theta)\mid\mathcal{F}_{k,h}],\quad\bar{g}_{k,h}(\theta)=\mathbb{E}_{y_{k,h}}[g_{k,h}(\theta)\mid\mathcal{F}_{k,h}].$
By Property 1, we have $f_{k,h}(\theta)$ is exponential concave such that
$f_{k,h}(\widehat{\theta}_{k,h})\leq f_{k,h}(\theta_{h}^{*})+\langle g_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$
Taking expectations on both sides, we have
$\bar{f}_{k,h}(\widehat{\theta}_{k,h})\leq\bar{f}_{k,h}(\theta_{h}^{*})+\langle\bar{g}_{k,h}(\theta_{h}^{*}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert(\widehat{\theta}_{k,h}-\theta_{h}^{*})\rVert_{W_{k,h}}.$ (14)
Based the property of KL diverge, we ensure the true parameter $\theta_{h}^{*}$ is the minimizer of the expected loss function by the following lemma.
Lemma 7.
For any $\theta\in\Theta$, we have $\bar{f}_{k,h}(\theta_{h}^{*})\leq\bar{f}_{k,h}(\theta)$.
Combining Lemma 7 and Equation (14), we have
00 0 $\leq\bar{f}_{k,h}(\widehat{\theta}_{k,h})-\bar{f}_{k,h}(\theta_{h}^{*})$
$\leq\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}$
$\leq\langle g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (15)
By standard analysis of OMD in Lemma 17, it holds
$2g_{k,h}(\widehat{\theta}_{k,h})(\widehat{\theta}_{k,h}-\theta_{h}^{*})\leq\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}}+\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}-\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}.$ (16)
Combining Equation (15) and Equation (16), we have
00 0 $\leq\frac{1}{2}\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+\frac{1}{2}\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\frac{1}{2}\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$
$\hskip 85.35826pt-\frac{\kappa}{2}\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle.$ (17)
First, we consider the first term . We show the gradient can be bounded by the following lemma.
Lemma 8.
For any positive semi-definite matrix $Z$, it holds that
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}\leq 4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{Z}^{2}.$ (18)
Then, we consider the second term. Note that $\widehat{\Sigma}_{k+1,h}=\widehat{\Sigma}_{k,h}+\frac{\kappa}{2}W_{k,h}$, we have
$\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}=\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+\frac{\kappa}{2}\lVert\widehat{\theta}_{k,h}-\theta^{*}\rVert_{W_{k,h}}^{2}.$ (19)
Next, we bound $\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$, which is a martingale difference sequence.
Lemma 9.
For any $\delta\in(0,1)$ and $\theta_{1},\ldots,\theta_{k},\theta^{*}\in[0,B]^{d}$, with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$\sum_{i=1}^{k}\langle\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i}),\theta_{i}-\theta_{h}^{*}\rangle\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$ (20)
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Combining (17), (18), and (19), we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}-\kappa\lVert\theta_{h}^{*}-\widehat{\theta}_{k,h}\rVert_{W_{k,h}}^{2}+\lVert g_{k,h}(\widehat{\theta}_{k,h})\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}$
$\hskip 28.45274pt+2\langle\bar{g}_{k,h}(\widehat{\theta}_{k,h})-g_{k,h}(\widehat{\theta}_{k,h}),\widehat{\theta}_{k,h}-\theta_{h}^{*}\rangle$
$\leq\lVert\widehat{\theta}_{k,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k,h}}^{2}+4\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$ (21)
where the first inequality holds by rearranging the terms in (17), the second holds by (18), (19) and (20). Summing (21) from $i=1$ to $k$, we have
$\lVert\widehat{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\widehat{\Sigma}_{k+1,h}}^{2}$ $\leq\lambda B^{2}+4\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{i+1,h}^{-1}}^{2}+(\frac{8}{\kappa}+16B)\gamma_{k}$
$\leq\lambda B^{2}+\frac{8}{\kappa}d\log\left(1+\frac{kU}{\lambda d}\right)+(\frac{8}{\kappa}+16B)\gamma_{k},$
where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, and the last inequality is by Lemma 6. This finishes the proof. ∎
B.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix B.1.
B.2.1 Proof of Lemma 7
Proof.
By the definition of $\bar{f}_{k,h}(\theta)$, we have
$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}\geq 0,$
where the last inequality is due to $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}$ is the Kullback-Leibler divergence between $p_{k,h}(\theta_{h}^{*})$ and $p_{k,h}(\theta)$, which always is non-negative. ∎
B.2.2 Proof of Lemma 8
Proof.
For any positive semi-definite matrix $Z$,
$(x_{i}-x_{j})^{\top}Z(x_{i}-x_{j})=x_{i}^{\top}Zx_{i}-x_{i}^{\top}Zx_{j}-x_{j}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq 0,\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}.$ (22)
which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}$.
Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
$\lVert g_{k,h}(\theta)\rVert_{Z}^{2}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$
$\hskip 56.9055pt+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})(p_{k,h}^{s^{\prime\prime}}(\theta)-y_{k,h}^{s^{\prime\prime}})(\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime\prime}}+\phi_{k,h}^{s^{\prime\prime}}Z\phi_{k,h}^{s^{\prime}})$
@@ B.2.2 Proof of Lemma 8
=\ = $2\sum_{s^{\prime}\in\mathcal{S}_{k,h}}(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})^{2}\phi_{k,h}^{s^{\prime}}Z\phi_{k,h}^{s^{\prime}}$

### v3

@@ Appendix B Properties of Multinomial Logit Function
Since $\lambda\geq 1$, we have $\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\leq 1$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$, we get
$\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\leq$ $2\sum_{i=1}^{k}\log\Big{(}1+\sum_{s^{\prime}\in\mathcal{S}_{i,h}}p_{i,h}(\theta_{h}^{*})\lVert\bar{\phi}_{i,h}^{s^{\prime}}\rVert^{2}_{\mathcal{H}_{i,h}^{-1}(\theta_{h}^{*})}\Big{)}$
[…]
where the last inequality holds by the determinant inequality in Lemma 14.
Proof of statement (II). The proof is same as that of (I), except that we replace $\theta_{h}^{*}$ with $\widetilde{\theta}_{i+1,h}$.
Proof of statement (III).  By Lemma 5, we have $H_{k,h}(\theta)\succeq\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}p_{k,h}^{s^{\prime}}(\theta)p_{k,h}^{\dot{s}_{k,h}}(\theta)\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. The remaining proof is the same as the proof of statement (I).
Proof of statement (IV).  By Lemma 5, we have $\forall\theta\in\Theta$, it holds that $\mathcal{H}_{k+1,h}(\theta)\succeq\mathcal{H}_{k,h}(\theta)+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\phi_{k,h}^{s^{\prime}}(\phi_{k,h}^{s^{\prime}})^{\top}$. Since $\lambda\geq 1$, we have $\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}(\theta)}\leq\kappa$. Using the fact that $z\leq 2\log(1+z)$ for any $z\in[0,1]$. By a similar analysis as the statement (I), we have
$\sum_{i=1}^{k}\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}^{2}$ $\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\kappa\max_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
$\leq\frac{2}{\kappa}\sum_{i=1}^{k}\log\Big{(}1+\kappa\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lVert\phi_{i,h}^{s^{\prime}}\rVert_{\mathcal{H}_{i,h}^{-1}}\Big{)}$
$\leq\frac{2}{\kappa}\log\left(\frac{\det(\mathcal{H}_{k+1,h}(\theta))}{\det(\mathcal{H}_{1,h}(\theta))}\right)$
$\leq\frac{2}{\kappa}d\log\left(1+\frac{k}{\lambda d}\right).$
This finishes the proof of statement (IV).
Proof of statement (V).  By (12), we have
Thus, we have
$\mathcal{H}_{k+1,h}\succeq\mathcal{H}_{k,h}+\kappa\sum_{s^{\prime}\in\dot{\mathcal{S}}_{k,h}}\widetilde{\phi}_{k,h}^{s^{\prime}}(\widetilde{\phi}_{k,h}^{s^{\prime}})^{\top}.$
Then, the remaining proof is similar to the proof of statement (III). ∎
Appendix C Useful Lemmas for MNL Mixture MDPs
In this section, we present some useful lemmas that are commonly used in the analysis.
C.1 Useful Lemmas
Lemma 7.
For any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$ and positive semi-definite matrix $\Lambda$, suppose $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$. Then, for any $V:\mathcal{S}\to[0,H]$ and $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{2})V(s^{\prime})}\bigg{|}\leq\epsilon_{s,a}^{\mathtt{1st}}+\epsilon_{s,a}^{\mathtt{2nd}}.$
where
@@ C.1 Useful Lemmas
$\epsilon_{s,a}^{\mathtt{1st}}$ $=H\beta\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})\Big{\lVert}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})\phi_{s,a}^{s^{\prime\prime}}\Big{\rVert}_{\Lambda^{-1}},$
$\epsilon_{s,a}^{\mathtt{2nd}}$ $=\frac{5}{2}H\beta^{2}\max_{s^{\prime}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\Lambda^{-1}}^{2}.$
Lemma 8.
Suppose $\forall(k,h,s,a)\in K\times[H]\times\mathcal{S}\times\mathcal{A}$ and $\widehat{\theta}_{k,h}\in\mathbb{R}^{d}$, it holds that $\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}$ where
$\widehat{\mathcal{C}}_{k,h}=\left\{\theta\ \Big{|}\ \bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)V(s^{\prime})}\bigg{|}\leq\Gamma_{k,h,s,a}\right\}.$ (13)
Define
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})\bigg{]}_{[0,H]},$ (14)
or,
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}\bigg{]}_{[0,H]},$ (15)
where $\widehat{V}_{k,h}(s)=\max_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s,a)$. Select the action as $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$. Then, for any $\delta\in(0,1]$, then it holds that
$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}\widehat{V}_{k,h+1}(s,a)+2\Gamma_{k,h,s,a}.$
Lemma 9.
Suppose Lemma 8 holds. Then, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{k,h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
C.2 Proof of Lemma 7
Proof.
By the second-order Taylor expansion at $\theta_{1}$, there exists $\bar{\theta}=\nu\theta_{1}+(1-\nu)\theta_{2}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{2})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})V(s^{\prime})}$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\theta_{1})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})+\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\theta_{2}-\theta_{1})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\theta_{2}-\theta_{1})V(s^{\prime})$
The gradient of $p_{s,a}^{s^{\prime}}(\theta)$ is given by
For the first-order term, we have
$\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\theta_{1})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})$
=\ = $\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{1})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})(\phi_{s,a}^{s^{\prime\prime}})^{\top}(\theta_{2}-\theta_{1})V(s^{\prime})$
[…]
=\ = $H\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}^{+}}p_{s,a}^{s^{\prime}}(\theta_{1})\bigg{(}\Big{(}\phi_{s,a}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{1})\phi_{s,a}^{s^{\prime\prime}}\Big{)}^{\top}\left(\theta_{2}-\theta_{1}\right)\bigg{)}$
@@ C.2 Proof of Lemma 7
where in the first inequality, we denote $\mathcal{S}_{h,s,a}^{+}$ as the subset of $\mathcal{S}_{h,s,a}$ such that $(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta_{2})(\phi_{s,a}^{s^{\prime\prime}})^{\top}(\theta_{2}-\theta_{1})$ is non-negative, the second inequality holds by the Holder’s inequality, and the third inequality is by the condition $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$.
For the second-order term, let $u_{s,a}^{s^{\prime}}(\theta)=(\phi_{s,a}^{s^{\prime}})^{\top}\theta$ and $p_{s,a}^{s^{\prime}}(u)=\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}}\exp(u_{s,a}^{s^{\prime\prime}})}$, further define
$F(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})},\quad\widetilde{F}(u)=\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\frac{\exp(u_{s,a}^{s^{\prime}})V(s^{\prime})}{1+\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\exp(u_{s,a}^{s^{\prime\prime}})}.$
Then, we have
$\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}(\theta_{2}-\theta_{1})^{\top}\nabla^{2}p_{s,a}^{s^{\prime}}(\bar{\theta})(\theta_{2}-\theta_{1})V(s^{\prime})$
=\ = $\frac{1}{2}\big{(}u(\theta_{2})-u(\theta_{1})\big{)}^{\top}\nabla^{2}\widetilde{F}(u(\bar{\theta}))\big{(}u(\theta_{2})-u(\theta_{1})\big{)}$
=\ = $\frac{1}{2}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}\big{(}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{)}^{\top}\frac{\partial^{2}\widetilde{F}(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{)}$
@@
where the inequality holds by $V(s)\in[0,H],\forall s$.
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$